The diagonals of a rhombus are along the line y=x and y=-x.


Area of rhombus square root 2 sq. units.


One vertex is (1,1), find other three vertices.

Grade:11

1 Answers

Chetan Mandayam Nayakar
13 years ago

Dear Karthik, 

the diagonals of a rhombus bisect each other perpendicularly. let A(1,1), B,C,D, be the four vertices. C is given by (-1,-1). let O be origin. (1/2)*AC*OD=√2/2, OD=1/2 units. Thus D is (-1/2√2,1/2√2), B is (1/2√2,-1/2√2)

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar respectively : Click here to download the toolbar.. 

Askiitians Expert

Chetan Mandayam Nayakar – IIT Madras